# e1: Learning Adaptive Control of Reasoning Effort

### v1

@@ Abstract
Increasing the thinking budget of AI models can significantly improve accuracy, but not all questions warrant the same amount of reasoning. Users may prefer to allocate different amounts of reasoning effort depending on how they value output quality versus latency and cost. To leverage this tradeoff effectively, users need fine-grained control over the amount of thinking used for a particular query, but few approaches enable such control. Existing methods require users to specify the absolute number of desired tokens, but this requires knowing the difficulty of the problem beforehand to appropriately set the token budget for a query. To address these issues, we propose Adaptive Effort Control, a self-adaptive reinforcement learning method that trains models to use a user-specified fraction of tokens relative to the current average chain-of-thought length for each query. This approach eliminates dataset- and phase-specific tuning while producing better cost-accuracy tradeoff curves compared to standard methods. Users can dynamically adjust the cost-accuracy trade-off through a continuous effort parameter specified at inference time. We observe that the model automatically learns to allocate resources proportionally to the task difficulty and, across model scales ranging from 1.5B to 32B parameters, our approach enables approximately 3x reduction in chain-of-thought length while maintaining or improving performance relative to the base model used for RL training.
@@ 1 Introduction
Additionally, across model scales from 1.5B to 32B parameters, our approach enables significant efficiency gains. On math tasks, we find that our training approach results in an approximately 3x reduction in chain-of-thought length while maintaining or even improving performance relative to the base model used for our RL training. Furthermore, after training, it is possible to calibrate the effort parameter to provide linear control over either relative accuracy or relative token usage, enabling intuitive user control tailored to different applications.
@@ 2 Related Work
Several works have shown that LLM performance can be improved by increasing computation at inference time in different ways. Methods can involve repeated sampling (Lightman et al., 2023; Brown et al., 2024), search-based techniques with a verifier (Wu et al., 2024; Snell et al., 2024; Manvi et al., 2024; Uscidda et al., 2025), or increasing the “reasoning,” or chain-of-thought computation, before producing an answer (Guo et al., 2025; Team et al., 2025). However, increased accuracy does not necessarily lead to more “intelligent” behavior Achille and Soatto (2025), as the models can simply improve accuracy by using the additional resources for extended brute force search, without learning new strategies. On the other hand, learning to jointly minimizing both inference time and error rate leads to maximizing the algorithmic information in the trained model (Achille and Soatto, 2025, Theorem 1.1).
@@ 5.1 Variable Effort Control and Improved Reasoning Efficiency
We first assess whether our approach enables variable effort control on math evaluation datasets. In Fig. 1a we show that after training, increasing the effort $r$ in the prompt leads to increasing accuracy and number of generated tokens across datasets of varying difficulty. We find that the model uses more tokens on more difficult problems, for example allocating more budget for problems in the AIME dataset which is generally considered more difficult than the Math500 dataset, where our model correctly allocates less budget. Note that in addition to allowing control, this training approach also allows us to significantly reduce the number of generated tokens while increasing accuracy (even allowing us to achieve over $60\%$ on AIME 2024 using a 7B model). Our training approach also leads to significant compression ($\sim$3x) compared to the base model used for RL training and models trained with standard reinforcement learning.
In Fig. 1b we show that our approach can be applied to models of various sizes (ranging from 1.5B to 32B), enabling more efficient reasoning and improved performance relative to the base model used for our RL training. We observe better performance vs number of generated token curves with increasing model size. We show aggregated results in Fig. 1b, and we show dataset-specific results in Appendix Fig. 13.
@@ 5.4 Calibration for interpretable control over percentage effort
When calibrating for the relative number of tokens, we define the maximum effort for a problem $x$ to be the average number of tokens generated $T(x,r=1)$ when using the largest value of $r$. On a small validation set (15 examples from each of Math500, AMC and AIME) we compute the desired percentage effort as $d(r)=\mathbb{E}_{x}\left[\frac{T(x,r)}{T(x,r=1)}\right]$, allowing us to map from user-specified desired percentage effort $d$ to $r$ (that is used to augment the prompt). On new examples, we similarly compute the outputted percentage effort as $o(x,r)=\frac{T(x,r)}{T(x,r=1)}$. We find that the outputted percentage effort is highly correlated with the desired percentage effort (Fig. 4b) even though the absolute number of tokens differs significantly across these datasets. We summarize these results by computing the average relative error $e=\mathbb{E}_{x\in D,r}[o(x,r)-d(r)]$ over the different datasets $D$, finding that the average relative error is $\sim$11% for each of these datasets (Fig. 14).
We similarly calibrate based on accuracy by computing $a(r)=\frac{\text{Acc(r)}}{Acc(r=1)}$ using the same validation set. We compare the desired accuracy $a(r)$ that a user can specify against the accuracy on the test set in Fig. 4a, finding an average error of $\sim$4%.
@@ 5.6 Effect of objective and robustness to prompt
Figure 7: Effect of RL training objective and robustness to prompt. We compare our effort objective a modified objective (Eq. 5) that penalizes the length of the generation regardless of problem difficulty. For this experiment we use a modified prompt that specifies the effort through a “points” parameter $p$. After training with the effort objective (circles in plot), we observe the ability to increase accuracy and the number of generated tokens by increasing $p$, highlighting that the dependence on the effort parameter does not depend on the specific prompt used. Training with a small length penalty ($\lambda=1e-4$) does not allow modulating the number of generated tokens or accuracy by varying the effort parameter $p$ (squares). Training with a small length penalty ($\lambda=1e-3$, triangles) leads to the ability to modulate the number of tokens and accuracy (within a narrow range), but results in significantly fewer generated tokens (less than 2000 for all datasets).
@@ 5.7 Selecting effort level for a task: Conceptual Depth of the trained model
[…]
$\kappa(\mathcal{D},T,\lambda_{\mathcal{D}})=E_{\mathcal{D}}(T)+\lambda_{\mathcal{D}}T,$ (6)
where $\lambda_{\mathcal{D}}$ represents the cost of time as determined by the environment, $E_{\mathcal{D}}$ denotes the error rate on a dataset $\mathcal{D}$, and $T$ denotes the average length of the chain-of-thought when solving queries from the task. We should emphasize that the above Eq. 6 is for evaluating a trained model, not for training a model to be an optimal solver (for that, see (Achille and Soatto, 2025)).
Figure 8: Conceptual Depth and task-specific effort selection. A user or agent should select the number of chain of thought tokens to balance the performance on the task and the cost incurred for achieving that performance. In math exams such as AMC (red) and AIME (green), the correct answer rendered after the time limit is worthless, and different exams have different time limits. When time and loss are compounded through the cost of time, the minimum value measures what we call the conceptual depth of the model. An “intelligent” agent should be able to choose this operating point, shown as black stars, depending on the specific task and query at hand. The error rate (where $\lambda=0$) is shown in a lighter shade. The optimal operating point that minimizes the conceptual depth can be specified by either the number of tokens or the effort level, and is approximately consistent across effort levels, even though the absolute number of tokens differs significantly depending on the particular exam and its time limit.
@@
$T^{*}(\mathcal{D},\lambda_{\mathcal{D}})=\arg\min_{T}\kappa(\mathcal{D},T,\lambda_{\mathcal{D}}).$
@@
In Fig. 8, we plot the conceptual depth for various values of $\lambda_{\mathcal{D}}$ in math tasks of varying difficulty. We set $\lambda_{\mathcal{D}}$ to be inversely proportional to the time constraints in actual testing scenarios: $\lambda_{MATH}=\lambda$, $\lambda_{AMC}=\lambda/3$, $\lambda_{AIME}=\lambda/12$, reflecting that students receive 12 minutes per AIME problem and 3 minutes per AMC problem, while we approximate (as this is not a standardized exam) 1 minute per MATH500 problem. We vary the value of $\lambda$ across the different panels while keeping the scaling across the datasets fixed. We show that even though the number of tokens needed to minimize $\kappa(\mathcal{D})$ varies significantly across the different math tasks, the effort level needed to minimize $\kappa(\mathcal{D})$ is similar across the different tasks for a given value of $\lambda$. We show this is the case for various values of $\lambda$ (shown across the 3 panels). For larger values of $\lambda$ the models use a smaller effort level, and for smaller values of $\lambda$, the models use a larger effort level.
The minimum of $\kappa(\mathcal{D})$ occurs at similar effort levels across the different math datasets as $\frac{dE_{\mathcal{D}}(T)}{dT}$ evaluated at the different effort levels is inversely proportional to the (real) time allocated per question, which scales $\lambda_{\mathcal{D}}$ (Fig. 19). Controlling the effort therefore provides a simple and practical knob to select an optimal operating point across different tasks with different cost-of-time.
@@ 6 Conclusion
We introduced AEC, a reinforcement learning method for enabling prompt-based control over the relative amount of chain-of-thought tokens for reasoning models, while being adaptive to the difficulty of a problem. In addition to enabling control over the generation length, our approach enables an approximately 33x reduction in chain of thought length while maintaining or improving performance relative to the base model.
@@ 7.3 Additional Figures
Figure 13: Same as Fig. 1a for various model sizes. These were aggregated to form Fig. 1c.
Figure 14: Calibration of percentage effort across datasets. We plot the average error (and standard deviation) across effort values from Fig. 4d for each dataset.

### v2

@@ Abstract
Increasing the thinking budget of AI models can significantly improve accuracy, but not all questions warrant the same amount of reasoning. Users may prefer to allocate different amounts of reasoning effort depending on how they value output quality versus latency and cost. To leverage this tradeoff effectively, users need fine-grained control over the amount of thinking used for a particular query, but few approaches enable such control. Existing methods require users to specify the absolute number of desired tokens, but this requires knowing the difficulty of the problem beforehand to appropriately set the token budget for a query. To address these issues, we propose Adaptive Effort Control, a self-adaptive reinforcement learning method that trains models to use a user-specified fraction of tokens relative to the current average chain-of-thought length for each query. This approach eliminates dataset- and phase-specific tuning while producing better cost-accuracy tradeoff curves compared to standard methods. Users can dynamically adjust the cost-accuracy trade-off through a continuous effort parameter specified at inference time. We observe that the model automatically learns to allocate resources proportionally to the task difficulty and, across model scales ranging from 1.5B to 32B parameters, our approach enables a 2-3x reduction in chain-of-thought length while maintaining or improving performance relative to the base model used for RL training.
@@ 1 Introduction
Additionally, across model scales from 1.5B to 32B parameters, our approach enables significant efficiency gains. On math tasks, we find that our training approach results in a 2-3x reduction in chain-of-thought length while maintaining or even improving performance relative to the base model used for our RL training. Furthermore, after training, it is possible to calibrate the effort parameter to provide linear control over either relative accuracy or relative token usage, enabling intuitive user control tailored to different applications.
@@ 2 Related Work
Several works have shown that LLM performance can be improved by increasing computation at inference time in different ways. Methods can involve repeated sampling (Lightman et al., 2023; Brown et al., 2024), search-based techniques with a verifier (Wu et al., 2024; Snell et al., 2024; Manvi et al., 2024; Uscidda et al., 2025), or increasing the “reasoning,” or chain-of-thought computation, before producing an answer (Guo et al., 2025; Team et al., 2025). However, increased accuracy does not necessarily lead to more “intelligent” behavior Achille and Soatto (2025), as the models can simply improve accuracy by using the additional resources for extended brute force search, without learning new strategies. On the other hand, learning to jointly minimizing both inference time and error rate leads to maximizing the algorithmic information in the trained model (Achille and Soatto, 2025).
@@ 5.1 Variable Effort Control and Improved Reasoning Efficiency
We first assess whether our approach enables variable effort control on math evaluation datasets. In Fig. 1a we show that after training, increasing the effort $r$ in the prompt leads to increasing accuracy and number of generated tokens across datasets of varying difficulty. We find that the model uses more tokens on more difficult problems, for example allocating more budget for problems in the AIME dataset which is generally considered more difficult than the Math500 dataset, where our model correctly allocates less budget. Note that in addition to allowing control, this training approach also allows us to significantly reduce the number of generated tokens while increasing accuracy (even allowing us to achieve over $60\%$ on AIME 2024 using a 7B model). Our training approach leads to significant reduction in the number of generated tokens compared to the base model used for RL training (3x reduction at equal performance, averaged over tasks) and models trained with standard reinforcement learning, while allowing for increasing performance with more generated tokens.
In Fig. 1b we show that our approach can be applied to models of various sizes (ranging from 1.5B to 32B), enabling more efficient reasoning while maintaining or improving performance relative to the base model used for our RL training. We observe better performance vs number of generated token curves with increasing model size. We show aggregated results in Fig. 1b, and we show dataset-specific results in Appendix Fig. 13. Across different math tasks and model sizes, we observe a 1.9-3.7x reduction in generated tokens when matching the baseline performance level, with larger reductions for smaller models (3.4x reduction averaged across tasks for 1.5B vs 2x for 32B; Fig. 13).
@@ 5.4 Calibration for interpretable control over percentage effort
When calibrating for the relative number of tokens, we define the maximum effort for a problem $x$ to be the average number of tokens generated $T(x,r=1)$ when using the largest value of $r$. On a small validation set (15 examples from each of Math500, AMC and AIME) we compute the desired percentage effort as $d(r)=\mathbb{E}_{x}\left[\frac{T(x,r)}{T(x,r=1)}\right]$, allowing us to map from user-specified desired percentage effort $d$ to $r$ (that is used to augment the prompt). On new examples, we similarly compute the outputted percentage effort as $o(x,r)=\frac{T(x,r)}{T(x,r=1)}$. We find that the outputted percentage effort is highly correlated with the desired percentage effort (Fig. 4b) even though the absolute number of tokens differs significantly across these datasets. We summarize these results by computing the average relative error $e=\mathbb{E}_{x\in D,r}[o(x,r)-d(r)]$ over the different datasets $D$, finding that the mean average relative error across datasets is 11% (Fig. 14).
We similarly calibrate based on accuracy by computing $a(r)=\frac{\text{Acc(r)}}{Acc(r=1)}$ using the same validation set. We compare the desired accuracy $a(r)$ that a user can specify against the accuracy on the test set in Fig. 4a, finding a mean absolute error (MAE) of 4% across effort levels and datasets.
@@ 5.6 Effect of objective and robustness to prompt
Figure 7: Effect of RL training objective and robustness to prompt. We compare our effort objective a modified objective (Eq. 5) that penalizes the length of the generation regardless of problem difficulty. For this experiment we use a modified prompt that specifies the effort through a “points” parameter $p$. After training with our AEC objective (circles), we observe the ability to increase accuracy and the number of generated tokens by increasing $p$, highlighting that the dependence on the effort parameter does not depend on the specific prompt used. Training with a small length penalty ($\lambda=1e-4$) does not allow modulating the number of generated tokens or accuracy by varying the effort parameter $p$ (squares). Training with a small length penalty ($\lambda=1e-3$, triangles) leads to the ability to modulate the number of tokens and accuracy (within a narrow range), but results in significantly fewer generated tokens (less than 2000 for all datasets).
@@ 5.7 Selecting effort level for a task: Conceptual Depth of the trained model
As we have anticipated, once the model is trained, its operation in a particular environment should be evaluated using a criterion that jointly accounts for space and time, with a unit conversion factor $\lambda$ that is specific to the environment, task, user, and model. We call the compound criterion that jointly accounts for space and time conceptual depth, which is a property of the trained model, and is related to various notions of complexity that also account for the time to produce the solution (e.g., Logical Depth (Bennett, 1988), Levin Complexity (Li and Vitanyi, 2019)). Results from Achille and Soatto (2025) suggest that, by incorporating time into training and optimizing our AEC objective, algorithmic information in the training data is transferred to the trained weights. The conceptual depth curve measures the tradeoff between error $E$ (or reward $R$) and time $T$ in a given environment, modulated by $\lambda_{\mathcal{D}}$:
$\kappa(T,\mathcal{D},\lambda_{\mathcal{D}})=E_{\mathcal{D}}(T)+\lambda_{\mathcal{D}}T,$ (6)
where $\lambda_{\mathcal{D}}$ represents the cost of time as determined by the environment, $E_{\mathcal{D}}$ denotes the error rate on a dataset $\mathcal{D}$, and $T$ denotes the average length of the chain-of-thought when solving queries from the task. For a given environment, the minimum value of this curve measures the conceptual depth of the model. We should emphasize that the above Eq. 6 is for evaluating a trained model, not for training a model to be an optimal solver (for that, see Achille and Soatto (2025)).
Figure 8: Conceptual Depth and task-specific effort selection. A user or agent should select the number of chain of thought tokens to balance the performance on the task and the cost incurred for achieving that performance. In math exams such as AMC (red) and AIME (green), the correct answer rendered after the time limit is worthless, and different exams have different time limits. When time and loss are compounded through the cost of time, the minimum value measures what we call the conceptual depth of the model. An “intelligent” agent should be able to choose this operating point, shown as black stars, depending on the specific task and query at hand. The error rate (where $\lambda=0$) is shown in a lighter shade. The optimal operating point can be specified by either the number of tokens or the effort level. For a given $\lambda$, the optimal effort level is similar across the different exams, even though the optimal number of tokens differs significantly depending on the particular exam and its time limit.
@@
$T^{*}(\mathcal{D},\lambda_{\mathcal{D}})=\arg\min_{T}\kappa(T,\mathcal{D},\lambda_{\mathcal{D}}).$
@@
In Fig. 8, we plot the conceptual depth curves for various values of $\lambda_{\mathcal{D}}$ in math tasks of varying difficulty. We set $\lambda_{\mathcal{D}}$ to be inversely proportional to the time constraints in actual testing scenarios: $\lambda_{MATH}=\lambda$, $\lambda_{AMC}=\lambda/3$, $\lambda_{AIME}=\lambda/12$, reflecting that students receive 12 minutes per AIME problem and 3 minutes per AMC problem, while we approximate (as this is not a standardized exam) 1 minute per MATH500 problem. We vary the value of $\lambda$ across the different panels while keeping the scaling across the datasets fixed. We show that even though the number of tokens needed to minimize $\kappa$ varies significantly across the different math tasks, the effort level needed to minimize $\kappa$ is similar across the different tasks for a given value of $\lambda$. We show this is the case for various values of $\lambda$ (shown across the 3 panels). For larger values of $\lambda$ the models use a smaller effort level, and for smaller values of $\lambda$, the models use a larger effort level.
The minimum of $\kappa$ occurs at similar effort levels across the different math datasets as $\frac{dE_{\mathcal{D}}(T)}{dT}$ evaluated at the different effort levels is inversely proportional to the (real) time allocated per question, which scales $\lambda_{\mathcal{D}}$ (Fig. 19). Controlling the effort therefore provides a simple and practical knob to select an optimal operating point across different tasks with different cost-of-time.
@@ 6 Conclusion
We introduced AEC, a reinforcement learning method for enabling prompt-based control over the relative amount of chain-of-thought tokens for reasoning models, while being adaptive to the difficulty of a problem. In addition to enabling control over the generation length, our approach enables a 2-33x reduction in chain of thought length while maintaining or improving performance relative to the base model.
@@ 7.3 Additional Figures
Figure 13: Same as Fig. 1a for various model sizes. These were aggregated to form Fig. 1b.
Figure 14: Calibration of percentage effort across datasets. We plot the average error (and standard deviation) across effort values from Fig. 4b for each dataset.